Question 4 < > Solve the initial value problem below. x+y - xy + y = 0, y(1) = – 5, y(1) = 0 y
Answer #1

given initial value problem x2y’’ - xy’ + y = 0 , y(1) = -5 and Y’(1) = 0

so this was in the form of   ax2y’’ + bxy’ + cy = 0 which was known as cauchy – euler homogeneous differential equation.

So the general solution is in the form of Y = Yc + Yp

Yc = homogeneous solution

Yp = particular solution

So for Yc we need to take homogeneous part   x2y’’ - xy’ + y = 0

So let us assume the general solution y = xm

Then y’ = m x m-1

And Y’’ = m(m-1) x m-2

Now substitute in given ivp x2y’’ - xy’ + y = 0

x2(m(m-1) x m-2) - x(mx m-1) + xm = 0

x2(m(m-1) x m x-2) - xmxmx-1 + xm = 0

m(m-1) x m - mxm + xm = 0

xm ([m(m-1) - m + 1])= 0

but xm is not equal to Zero so [m(m-1) - m + 1]= 0

m2-m - m + 1= 0 => m(m-1) – (m-1)= 0 => (m-1)(m-1) = 0

so m1 = m2 = 1

So when the roots are real and repeat then the general solution is in the form of

Y(x) = C1xm1 + C2 xm2 (lnx)

Y(x) = C1x + C2 x (lnx)

Now we have initial condition y(1) = -5

Y(x) = C1x + C2 x (lnx)

Y(1) = C1(1) + C2 (1) (ln(1))

-5 = C1(1) + C2 (1) (0)

-5 = C1

Now we have y’(1) = 0

Y(x) = C1x + C2 x (lnx)

Y’(x) = C1 + C2 (lnx) + C2 x (1/x)

Y’(x) = C1 + C2 (lnx) + C2

Y’(1) = C1 + C2 (ln(1)) + C2

0 = C1 + C2 (0) + C2

0 = C1 + C2 => C2 = -C1

And C1 = -5 then C2 = -(-5) => C2 = 5

Now the general solution

Y(x) = -5x + 5 x (lnx)

Y(x) = ((lnx) – 1) 5x